OpenStudy (anonymous):

Change to standard circle form: 3x^2+3y^2+5x-4y=1

OpenStudy (anonymous):

complete the squares for x and y. the remainder + 1 = radius

OpenStudy (anonymous):

I don't know how to complete the square with coefficients attached to the variables with the 2nd degree.

OpenStudy (anonymous):

you can look at it as 3(x^2 + 5x/3) + 3(y^2 - 4y/3) = 1 and complete the square inside the brackets :)

OpenStudy (anonymous):

Would I need to multiple whatever I add inside by 3 and add it to the other side?

OpenStudy (anonymous):

no. you can leave the 3 outside and forget about it. your answer will look like: 3(x + a/b)^2 + 3(y - c/d)^2 = e

OpenStudy (anonymous):

Why wouldn't it be (x^2 + 5x/3) + (y^2 - 4y/3) = 1/3

OpenStudy (anonymous):

you an also divide e by 3 to get rid of the 3's, in this case

OpenStudy (anonymous):

merp

OpenStudy (anonymous):

only because the coefficients are the same; rarely happens

OpenStudy (anonymous):

(x+5/6)^2+(y-2/3)^2=53/36 Good or no good?